solve boy substitution (8x +2y=4 ca 3x - 5y = 11
Answer #1

Griven that is 8x+2y= 40 x-59 = 11 20x5+2x9 yox Hop=20 2x-104 = 22 + 42x=42 x=11 X=1 in 80 +2y=4 24-u-8 24 =-4 y=-2 x=1 y=-2